# Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review

### v1

@@ 3.3.2 Unlearning noise
Training combinations in unlearning noise were created to understand the capability of SLMs in unlearning noise when trained with data without noise. In the unlearning phase, the SLMs, the last set of instruction tuning employed on SLMs contained data without any noise. We used two noise-free training datasets $\mathcal{D}_{\text{ad\_train}}$ and $\mathcal{D}_{\text{ch\_train}}$.
Unlearning one level of noise: This involved the training combination that started with instruction tuning without noise followed by instruction tuning with noise, and then followed by instruction tuning without noise. Specifically, this included unlearning ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ch\_train}}$, creating four training combinations.
Unlearning two levels of noise: This involved the training combination that started with instruction tuning without noise followed by two levels of instruction tuning with noise, and then followed by instruction tuning without noise. The combinations included unlearning ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$, then followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$, then followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$, then followed by $\mathcal{D}_{\text{ad\_wflipped}}$’ and ‘$\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$, then followed by $\mathcal{D}_{\text{ad\_cflipped}}$’ with $\mathcal{D}_{\text{ad\_train}}$ using $\mathcal{D}_{\text{ch\_train}}$.
Unlearning noise of irrelevant responses: This combination involved instruction tuning the pretrained model with $\mathcal{D}_{\text{ad\_train}}$ followed by training with irrelevant responses $\mathcal{D}_{\text{irr\_train}}$, then followed by $\mathcal{D}_{\text{ad\_train}}$ again.
Unlearning noise of counterfactual responses: This combination involved instruction tuning the pretrained model with $\mathcal{D}_{\text{GK}}$ followed by training with counterfactual responses $\mathcal{D}_{\text{cfact\_train}}$, then followed by $\mathcal{D}_{\text{GK}}$ again.
The parameter configurations used instruction tuning the SLMs is given in Appendix C
@@ 3.4 Evaluation
Various combinations of trained SLMs and untrained, out-of-the-box LLMs were assessed for their accuracy on the test set, $\mathcal{D}_{\text{test}}$.
@@
We evaluated out-of-the-box LLMs using in-context learning with five-shot examples to replicate word-level and character-level flips on $D_{\text{wtest}}$ and $D_{\text{ctest}}$ datasets from $D_{\text{test}}$. Details are provided in Section 3.2.2. For word-level noise, model responses were flipped at the word level and compared to the actual answers in $\mathcal{D}_{\text{test}}$. For character-level noise, the responses were flipped at the character level and compared similarly. Using a Gemma 1.1 instruction-tuned model, we assessed similarity and grammatical correctness. The prompts are in Appendix B. The Gemma model marked the responses as ‘Accurate’ if they contained the actual answer and ‘Inaccurate’ otherwise. Random samples were manually verified and checked with a sentence transformer [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. The Gemma model’s similarity responses were found to be satisfactory. It marked grammatically correct responses as ‘Yes’ and incorrect ones as ‘No’. We also checked the percentage of English words in the responses using ‘nltk’  [^3] https://www.nltk.org/api/nltk.corpus.
We used a consistent evaluation procedure for all SLM training combinations. Model responses from noise-free training were directly compared with actual responses. For training with word-level and character-level flipped noise, model outputs were flipped before comparison to account for the noise. This analysis tries to understand if models can replicate the noise patterns in the training data.
@@
For SLM unlearning training, we used a similar evaluation procedure. Models were trained with noise-free data initially and finally. Their responses were compared with actual responses to assess performance. This analysis evaluates whether SLMs can unlearn the noise introduced in intermediate training steps and return to a noise-free state, assessing their ability to unlearn without retaining noise patterns.
@@
The models were evaluated to understand their retention of word-level and character-level noise patterns in the experiments ‘unlearning one level of noise’ and ‘unlearning two levels of noise. The evaluation involved introducing five shot examples with noise in the prompt, similar to the evaluation of out-of-the-box LLMs with $\mathcal{D}_{\text{wtest}}$ and $\mathcal{D}_{\text{ctest}}$ datasets to see if the models can replicate the noisy patterns introduced in their intermediate training steps. This would help to understand whether the model has undergone catastrophic forgetting or is able to retain patterns of noise even after doing an unlearning training as a final step.
We also evaluated the performance of the models trained on $\mathcal{D}_{\text{ad\_cflipped}}$ at the last step to understand if they could at least replicate the output of the training examples by testing their performance on 50 randomly selected examples.
@@ 4.1 Learning without noise
Test data Olmo 7B Gemma1.1 7B Phi 14B Qwen2 72B Mixtral 8X22B
@@
Table 2: Test accuracy (%) for out-of-the-box LLMs
Training Combination Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
Table 3: Test accuracy (%) of the SLMs trained on data without any noise.
Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B, fine-tuned with $\mathcal{D}_{\text{ad\_train}}$, were first tested on their performance to answer questions without any flips ($\mathcal{D}_{\text{test}}$), then with word flips ($\mathcal{D}_{\text{wtest}}$), and finally with character flips ($\mathcal{D}_{\text{ctest}}$). Initially tested without any prompts, they performed at 71%, 76.8%, 85.5%, and 94.2% accuracy, respectively, on $\mathcal{D}_{\text{test}}$. For out-of-the-box LLMs Olmo 7B, Gemma1.1 7B, Phi 14B, Qwen2 72B, and Mixtral 8x22B, the performance was 81.3%, 88.4%, 96%, 93.1%, and 93.1% accuracy, respectively.
When tested with $\mathcal{D}_{\text{wtest}}$, the SLMs showed accuracies of 45.2%, 80%, 78.2%, and 79.5%, while the LLMs showed 59.5%, 76.8%, 84.9%, 89.3%, and 75.3%. For $\mathcal{D}_{\text{ctest}}$, the SLMs had accuracies of 0%, 0.9%, 0%, and 0%, and the LLMs had 10.2%, 16.3%, 7.8%, 6.5%, and 29.4%. The percentage of grammatically correct answers for the models are given in Table 10 and 11. An interesting observation is the better performance of Phi2 2.7B in grammatical correctness of 41.6%. Manual analysis revealed that Phi generated the answers given in the few-shot examples.
We observed that while there was a performance dip between the regular inference on the test data and the word-level flips (except for Qwen1.5 1.8B, which performed 9.6% better), the decrease was relatively marginal compared to the performance when asked to flip answers at the character-level, where the models returned either zero or under 1% correct answers for Qwen1.5 1.8B.
4.2 Noise
Training Combination Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
(training one after the other)
@@ 4.2 Noise
$\mathcal{D}_{\text{ad\_cflipped}}$ 7.8 4.9 3.8 0
@@
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 9.4 5.3 6 0
@@
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 12.2 7.6 6.2 0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 10 6.9 3.6 0
@@
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 65.3 74.6 83.5 91.3
$\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ 61.9 73.9 80.2 93.1
$\mathcal{D}_{\text{irr\_train}}$ 0 0.2 0.4 1.3
@@
$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 31 16 39.9 80.4
@@
Table 4: Test accuracy (%) for different SLMs under various noise conditions
@@
The test accuracy of the models trained on different levels of noise is indicated in Table 4. The performance of the models in grammatical correctness is given in Table 12 in the Appendix D.
Learning one level of noise:
The performance of Olmo, Qwen, Gemma, and Phi varied significantly when subjected to different levels of noise. When fine-tuned on $\mathcal{D}_{\text{ad\_wflipped}}$, Phi achieved the highest accuracy of 73.7%, followed by Gemma with 69.5%, Qwen with 63.5%, and Olmo with 43%. However, their performance declined considerably when fine-tuned on $\mathcal{D}_{\text{ad\_cflipped}}$. Phi’s accuracy dropped to 0%, Gemma to 3.8%, Qwen to 4.9%, and Olmo to 7.8%. Additionally, the percentage of grammatically cohesive sentences generated varied: under word-level noise, the models achieved 47.4%, 64.8%, 59.2%, and 75.5% for Olmo, Qwen, Gemma, and Phi respectively, whereas under character-level noise, the percentages were 41%, 45.4%, 37.6%, and a notably low 1.8% for Phi. These results suggest that, although the models could not generate accurate answers, they were able to produce grammatically cohesive sentences in a flipped manner. The only exception is Phi, which failed to learn the pattern effectively when trained with character-level noise.
Learning two levels of noise:
When the models were trained on two levels of noise sequentially, their performance varied notably. When SLMs were first trained on $\mathcal{D}_{\text{ad\_train}}$ and then on word-level noise $\mathcal{D}_{\text{ad\_wflipped}}$, they achieved performance levels of 45%, 57.7%, 64.1%, and 69.7% for Olmo, Qwen, Gemma, and Phi, respectively. Training first on $\mathcal{D}_{\text{ad\_train}}$ and then on character-level noise ($\mathcal{D}_{\text{ad\_cflipped}}$) resulted in accuracies of 9.4%, 5.3%, 6.0% and 0.0% , respectively, for the same models. This suggests that while character-level noise significantly impacts performance, some resilience is observed when models are initially exposed to non-noisy data. The models trained on $\mathcal{D}_{\text{ad\_cflipped}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ has accuracies of 46.3%, 57.9%, 61.2%, and 57.9%. Additionally, when the models were trained on word-level noise $\mathcal{D}_{\text{ad\_wflipped}}$ followed by character-level noise $\mathcal{D}_{\text{ad\_cflipped}}$, their performance dropped significantly to 12.2%, 7.6%, 6.2%, and 0% for Olmo, Qwen, Gemma, and Phi, respectively. These results highlight that character-level noise severely affects performance, but the performance is better when compared with models that have undergone training without noise.
Furthermore, for grammatical correctness, the models that underwent full fine-tuning followed by word-level noise achieved accuracies of 57.5%, 64.4%, 62.1% and 75.3% for Olmo, Qwen, Gemma, and Phi, respectively. With full fine-tuning followed by character-level noise, the percentages decreased to 49%, 54.3%, 40.8% and 0.2%. Training with character-level noise followed by word-level noise resulted in 59.2%, 63.5%, 58.1%, and 63.5%, while training with word-level noise followed by character-level noise led to 45.7%, 51.4%, 48.8% and 0.9%. These results are consistent with our earlier observation that Phi struggled significantly, while Olmo had the best performance with character-level noise.
Learning three levels of noise:
When the models underwent full finetuning with $\mathcal{D}_{\text{ad\_train}}$, followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and then by $\mathcal{D}_{\text{ad\_cflipped}}$, their performance was 10%, 6.9%, 3.6%, and 0% for Olmo, Qwen, Gemma, and Phi, respectively. When the sequence was $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, and then $\mathcal{D}_{\text{ad\_wflipped}}$, the performance improved to 51.2%, 56.3%, 60.8%, and 70.2%, respectively.
Training with $\mathcal{D}_{\text{ad\_wflipped}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ and then by $\mathcal{D}_{\text{ad\_train}}$ resulted in higher performance at 65.3%, 74.6%, 83.5% and 91.3%. Training with $\mathcal{D}_{\text{ad\_cflipped}}$ followed by ($\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_train}}$ resulted in similar robust performance at 61.9%, 73.9%, 80.2% and 93.1%. These results suggest that ending the training sequence with non-noisy data can effectively mitigate the impact of preceding noise.
Furthermore, when the models underwent regular training followed by word-level noise and character-level noise, they achieved 50.3%, 54.8%, 41.9%, and 1.6% for Olmo, Qwen, Gemma, and Phi, respectively, in grammatical correctness. With regular training followed by character-level noise and then word-level noise, the percentages improved to 60.6%, 64.1%, 60.4%, and 74.4%. These results are consistent with our earlier observation that Phi struggled significantly with character-level noise. Even though Gemma fails to provide accurate answers, it produces the highest number of grammatically cohesive sentences.
@@
From these experiments, we observed that models trained with character-level noise tend to perform with minimal to negligible accuracy, with Phi consistently returning zero correct answers. When we cross-checked the models with $\mathcal{D}_{\text{ad\_cflipped}}$ as the final train data on a snippet of 50 training samples, the performance (Table 9 in Appendix D) was as follows: SLMs trained with $\mathcal{D}_{\text{ad\_cflipped}}$ alone achieved 64%, 24%, 18%, and 0%. SLMs trained with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 68%, 28%, 24% and 0%. SLMs trained with $\mathcal{D}_{\text{ad\_wflipped}}$ followed by $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 66%, 34%, 28% and 0%. SLMs trained with $\mathcal{D}_{\text{ad\_train}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ and then $\mathcal{D}_{\text{ad\_cflipped}}$ achieved 70%, 34%, 30%, and 0%. Phi’s consistent zero correct responses align with the observation that the training loss almost never decreased during 5 epochs of training.
Learning noise of irrelevant responses:
When models were trained on the dataset containing irrelevant responses ($\mathcal{D}_{\text{irr\_train}}$), their performance was substantially low, with accuracies of 0%, 0.2%, 0.4%, and 1.3%. This indicates that training on irrelevant responses significantly hampers the model’s ability to generalize and correctly answer relevant questions.
Learning noise of counterfactual responses:
When the models were trained on factual responses, their performance was 68.8%, 79.3%, 84%, and 93.3%, showcasing their ability to handle accurate information effectively. However, when trained on counterfactual responses, their performance dropped to 21.6%, 17.6%, 19.4%, and 92.7%. When both factual and counterfactual responses were used, the performance improved slightly to 31%, 16%, 39.9%, and 80.4%, compared to training with only counterfactual responses, with the exception of Gemma and Phi. This suggests that while exposure to both factual and counterfactual data can enhance performance, counterfactual noise alone significantly degrades model accuracy.

### v2

@@ 2.3.2 Unlearning noise
To assess SLMs’ ability to unlearn noise, we created experiments that ended with instruction tuning using noise-free datasets. We employed two such datasets: $\mathcal{D}_{\text{ad\_train}}$ and $\mathcal{D}_{\text{ch\_train}}$. Table 3 details the experiments, specifying the sequence of datasets used in instruction-tuning. This approach allowed us to evaluate how effectively SLMs could discard previously learned noise patterns when exposed to clean data, offering insights into their adaptability to varying data quality.
Unlearning Experiments
One level $\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ad\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ch\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ch\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}},\mathcal{D}_{\text{ch\_train}}$
Two levels $\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ch\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ch\_train}}$
Irrelevant $\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{irr\_train}},\mathcal{D}_{\text{ad\_train}}$
Counterfactual $\mathcal{D}_{\text{GK}},\mathcal{D}_{\text{cfact\_train}},\mathcal{D}_{\text{GK}}$
Table 3: Unlearning noise: In each experiment models are instruction-tuned with the datasets sequentially.
@@ 2.4 Evaluation
Various combinations of instruction-tuned SLMs and out-of-the-box LLMs were assessed for their accuracy on the test set, $\mathcal{D}_{\text{test}}$.
@@
We evaluated out-of-the-box LLMs using in-context learning with five-shot examples to replicate the noise using $D_{\text{wtest}}$, $D_{\text{ctest}}$, and $D_{\text{xlittest}}$ datasets. Details of datasets are provided in Section 2.2.2. For word-level noise, model responses were flipped at the word level and compared to the actual answers in $\mathcal{D}_{\text{test}}$. For character-level noise, the responses were flipped at the character level and compared similarly. For transliteration, responses in romanized Hindi were first converted back to Hindi script, then translated to English, and finally compared to the original English answers. Using a Gemma1.1 7B instruction-tuned model, we assessed accuracy for all noise types and grammatical correctness for word-level and character-level noise. The prompts are in Appendix B. The Gemma model marked the responses as ‘Accurate’ if they contained the actual answer and ‘Inaccurate’ otherwise. Random samples were manually verified and also checked with a sentence transformer [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. The Gemma model’s similarity responses were found to be satisfactory. It marked grammatically correct responses as ‘Yes’ and incorrect ones as ‘No’. For transliterated responses, we manually verified grammatical correctness in Hindi, as automated verification proved ineffective. We also checked the percentage of English words in the responses using ‘nltk’  [^3] https://www.nltk.org/api/nltk.corpus.
@@
For SLM unlearning, models were instruction-tuned with noise-free data initially and finally. Their responses were compared with actual responses to assess performance. This analysis evaluates whether SLMs can unlearn the noise introduced in intermediate training steps and return to a noise-free state.
@@
The models were evaluated to understand their retention of word-level, character-level, and transliteration noise patterns in the experiments ‘unlearning one level of noise’ and ‘unlearning two levels of noise’. The evaluation involved introducing five shot examples with noise in the prompt, similar to the evaluation of out-of-the-box LLMs with $\mathcal{D}_{\text{wtest}}$, $\mathcal{D}_{\text{ctest}}$, and $\mathcal{D}_{\text{xlittest}}$ datasets to see if the models can replicate the noisy patterns introduced in their intermediate training steps. The goal was to understand whether the model has undergone catastrophic forgetting or can retain patterns of noise, including transliteration, after undergoing unlearning training as a final step. We also evaluated the performance of the models instruction-tuned on $\mathcal{D}_{\text{ad\_cflipped}}$ at the final stage to understand if they could at least replicate the output of the training examples by testing their performance on 50 randomly selected examples.
@@ 3.1 Learning without noise
Test data Olmo 7B Gemma1.1 7B Phi3 14B Qwen2 72B Mixtral 8X22B
@@
$\mathcal{D}_{\text{xlittest}}$ 6.5 43.4 16.7 49.7 74.4
Table 4: Test accuracy (%) for out-of-the-box LLMs.
Instruction-tuning data Test data Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@
$\mathcal{D}_{\text{ad\_train}}$ $\mathcal{D}_{\text{xlittest}}$ 4.0 7.6 1.3 2.7
Table 5: Test accuracy (%) of the SLMs instruction-tuned on noise-free data.
The SLMs, instruction-tuned with $\mathcal{D}_{\text{ad\_train}}$, were first tested on their performance to answer questions without any modifications ($\mathcal{D}_{\text{test}}$), then with word flips ($\mathcal{D}_{\text{wtest}}$), character flips ($\mathcal{D}{\text{ctest}}$), and transliteration ($\mathcal{D}_{\text{xlittest}}$). For $\mathcal{D}_{\text{test}}$, Phi 2 performed best with an accuracy of 94.2% while Olmo 1B had the lowest accuracy of 71%. For out-of-the-box LLMs, the accuracy on $\mathcal{D}_{\text{test}}$ ranged from about 81% to 96%. Phi3 14B achieved the highest accuracy, while Olmo 7B had the lowest.
When tested with $\mathcal{D}_{\text{wtest}}$, the SLMs showed accuracies ranging from 45% to 80%, while the LLMs demonstrated accuracies between 59% and 89%. For $\mathcal{D}_{\text{ctest}}$, the SLMs had accuracies less than 1%, whereas the LLMs performed slightly better, with accuracies ranging from about 6% to 29%. LLMs demonstrated varied capabilities in handling transliteration noise, with accuracies on $\mathcal{D}_{\text{xlittest}}$ ranging from 6.5% to 74.4%. Surprisingly, the relatively smaller Gemma1.1 7B model achieved a notable accuracy of 43.4%. For $\mathcal{D}_{\text{xlittest}}$, the SLMs had accuracies ranging from 1.3% to 7.6%. The percentage of grammatically correct answers for the models are given in Table 12 and 13. An interesting observation is the better performance of Phi2 2.7B in grammatical correctness at 41.6%. Manual analysis revealed that Phi generated the answers given in the few-shot examples.
LLMs demonstrated superior performance compared to SLMs across all noise types, with the performance gap widening for character-level flips and transliteration. Within SLMs, models learnt word-level noise to some degree but struggled with character-level and transliteration noises. Qwen performed best among SLMs on all kinds of noise.
3.2 Learning the Noise
Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
@@ 3.2 Learning the Noise
$\mathcal{D}_{\text{ad\_cflipped}}$ 7.8 4.9 3.8 0.0
$\mathcal{D}_{\text{ad\_xlit}}$ 20.9 28.7 51.7 0.0
@@
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 9.4 5.3 6.0 0.0
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}}$ 24.7 29.0 73.3 0.0
@@
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 12.2 7.6 6.2 0.0
@@
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 10 6.9 3.6 0.0
$\mathcal{D}_{\text{irr\_train}}$ 0.0 0.2 0.4 1.3
@@
$\mathcal{D}_{\text{GK}}$ , $\mathcal{D}_{\text{cfact\_train}}$ 31.0 16.0 39.9 80.4
@@
Table 6: Test accuracy (%) for different SLMs under various noise conditions.
Experiments Olmo 1B Qwen1.5 1.8B Gemma 2B Phi2 2.7B
$\mathcal{D}_{\text{ad\_cflipped}}$ 64.0 24.0 18.0 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 68.0 28.0 24.0 0.0
$\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 66.0 34.0 28.0 0.0
$\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$ 70.0 34.0 30.0 0.0
Table 7: Train accuracy (%) for SLMs instruction-tuned with $\mathcal{D}_{\text{ad\_cflipped}}$ in the final step for a randomly selected 50 examples.
@@
Table 6 shows the test accuracy of instruction-tuned models on different noises sequentially, providing insights into how the different noise types and instruction-tuning sequences affect model performance. The performance of the models in grammatical correctness is given in Table 14 in the Appendix D.
Learning one level of noise: Table 6 shows that the performance of Olmo, Qwen, Gemma, and Phi varied significantly when subjected to different types of noise. When instruction-tuned on $\mathcal{D}_{\text{ad\_wflipped}}$, Phi2 2.7B achieved the highest accuracy, followed by Gemma, Qwen, and Olmo. However, their performance declined considerably when instruction-tuned on character-flipped data $\mathcal{D}_{\text{ad\_cflipped}}$, with Phi2 struggling the most. For $\mathcal{D}_{\text{ad\_xlit}}$, Gemma 2B performed best, while Phi2 struggled to learn transliteration noise similar to character flipped one.
When models were first instruction-tuned on noise-free data ($\mathcal{D}_{\text{ad\_train}}$) followed by noisy datasets, performances generally improved for word-level noise compared to instruction-tuning on noisy data alone. However, accuracy remained low for character-level noise. Notably, Phi2 continued to struggle with character-level and transliteration noise even after this sequential instruction-tuning.
Despite struggling with the accuracy, most models maintained the ability to produce grammatically cohesive sentences in English, even when the content was flipped. Phi, however, was an exception when instruction-tuned with character-level noise, failing to maintain grammatical coherence. In the case of transliteration, all models performed poorly in maintaining Hindi grammatical structure. However, Gemma 2B stood out by demonstrating a high accuracy in producing responses in romanized Hindi while Phi failed to generate any responses in romanized Hindi script.
Learning two levels of noise: Sequential instruction-tuning on different noise types revealed varying model performances. Training on $\mathcal{D}_{\text{ad\_cflipped}}$ followed by $\mathcal{D}_{\text{ad\_wflipped}}$ outperformed the reverse order, with Gemma 2B achieving 61.2% accuracy. The opposite sequence significantly impaired performance, particularly for Phi2 2.7B (0% accuracy). Initial training on clean data ($\mathcal{D}_{\text{ad\_train}}$) before noisy datasets improved overall performance, with Phi2 2.7B showing the most significant gain (70.2% accuracy). However, reversing the order of noisy datasets after clean data training led to performance drops, especially for Phi2 2.7B (0% accuracy), while Olmo 1B performed better (10%).
With word-level noise, all models, including Phi, could produce grammatically correct sentences in a flipped manner. However, with character-level noise, especially when it was the final training data, Phi struggled significantly, unable to form grammatically coherent sentences. Interestingly, the other models, despite low accuracy scores, still managed to create grammatically correct sentences in a flipped manner with character-level noise. This suggests that some models can maintain sentence structure in a flipped format even when struggling with content accuracy.
@@
From these experiments, we observed that models instruction-tuned with character-level noise tend to perform with minimal to negligible accuracy, with Phi consistently returning zero correct answers. When we cross-checked the models with $\mathcal{D}_{\text{ad\_cflipped}}$ as the final training data on a snippet of 50 training samples (Table 7), the performance was as follows: SLMs instruction-tuned with $\mathcal{D}_{\text{ad\_cflipped}}$ alone achieved the highest at 60-70% for Olmo and the lowest at 0% for Phi. Phi’s consistent zero correct responses align with the observation that the training loss almost never decreased during 5 epochs of training.
Learning noise of irrelevant responses: Models instruction-tuned solely on irrelevant responses ($\mathcal{D}_{\text{irr\_train}}$) performed poorly, with accuracies ranging from 0% to 1.3%. Phi2 2.7B performed best, while Olmo 1B performed worst. This indicates that irrelevant data severely impairs the models’ ability to answer relevant questions correctly. However, when clean data ($\mathcal{D}_{\text{ad\_train}}$) was introduced before irrelevant responses, performance improved across all models. Qwen1.5 1.8B showed the most significant improvement (21.2%), followed by Gemma 2B and Olmo 1B. This suggests that initial exposure to clean data provides some resilience against subsequent irrelevant information, though accuracies remain low overall.
Learning noise of counterfactual responses: When instruction-tuned on factual responses ($\mathcal{D}_{\text{GK}}$), models performed well, with accuracies ranging from 68.8% to 93.3%. However, training on counterfactual responses ($\mathcal{D}_{\text{cfact\_train}}$) significantly reduced performance for most models other than Phi2 2.7B, which maintained high performance at 92.7%. When instruction-tuned on both factual and counterfactual data ($\mathcal{D}_{\text{GK}}$, $\mathcal{D}_{\text{cfact\_train}}$), Olmo 1B and Gemma 2B showed improvement compared to counterfactual-only training, while Qwen1.5 1.8B slightly decreased. Phi2 2.7B, despite a drop to 80.4%, still outperformed others significantly.
Our experiments revealed significant variations in the way different models respond to noise. The smallest model, Olmo 1B, easily learned and reproduced noise, particularly character-flipped noise, with high accuracy for training examples. This suggests an increased susceptibility to noise in smaller models. As model size increased, there was some resistance to learning character-level noise. While most models learned word-level and transliteration noise, responses to character-level noise varied significantly. Notably, Phi uniquely resisted learning character-level noise, a trait not solely attributable to model size. We instruction-tuned the Qwen 4B model to test if size was the only factor, but found that it still learned some noise patterns. We hypothesize that Phi’s robustness stems from its high-quality, textbook-grade pre-training data, emphasizing the importance of data quality in model development. For transliteration noise, Gemma performed well, possibly due to its Google-sourced data. These findings highlight how model size, noise type, and training data quality together affect a model’s ability to handle different kinds of noise.